# CADSpotting: Robust Panoptic Symbol Spotting on Large-Scale CAD Drawings

### v3

@@ 4.3 Quantitative Evaluation
To demonstrate the effectiveness of our approach, we compare it against SOTA methods on the FloorPlanCAD [5] and LS-CAD datasets. A detailed comparison of semantic, instance, and panoptic symbol spotting tasks is provided in the following paragraphs.
@@
Semantic Symbol Spotting. As shown in Table.2, we compare our method with PanCADNet [5], CADTransformer [6], GAT-CADNet [40], and SymPoint [16]. Our CADSpotting approach achieves the highest performance in both the F1 and wF1 metrics, outperforming SymPoint by 13.6% in mIoU.
@@
Instance Symbol Spotting. We also evaluate our method against DINO [38] and SymPoint for instance symbol spotting, using the bounding box maximization of predicted mask to compute the AP metric, consistent with SymPoint. As shown in Table.3, our method achieves superior results in the AP50, AP75 and mAP metrics, with improvements of 5.9% AP50, 8.4% AP75 and 8.9% mAP over SymPoint.
@@
Panoptic Symbol Spotting. We then compare our method with the same methods used in the semantic symbol spotting comparison. The results, presented in Table.4, show that our approach surpasses SymPoint and other prior methods across all metrics, including Total/Thing/Stuff PQ, SQ and RQ. Additionally, our approach offers superior generalization across diverse CAD primitives, thanks to its dense point sampling based feature representation.
We also compare the effectiveness of Sliding Window Aggregation (SWA) and Block Partitioning (BP) methods when applied to the CADSpotting model on the LS-CAD dataset. We apply BP method using the same size as drawings the FloorPlanCAD dataset without overlapping between blocks, to divide large-scale CAD drawings into smaller blocks. When employing the BP method, we ensure that each primitive is uniquely represented by retaining only those whose starting points fall within the block. The results in Table 5 indicate that, compared to the BP method, SWA achieves the highest performance on PQ, with an absolute improvement of 10.4%, although there is a slight decrease of 1.0% on SQ. The result shows that our CADSpotting method with the SWA technique can accurately recognize instances of CAD symbols in large-scale CAD drawings.
4.4 Qualitative Comparison
@@
| Ours | PTv3 | 88.9 | 95.6 | 93.0 |
| Method | PQ | SQ | RQ |
| w/o Color Feature | 85.8 | 92.5 | 92.7 |
| Color Feature | 88.9 | 95.6 | 93.0 |
4.5 Ablation Studies
In this section, we present ablation studies examining the effects of feature learning and RGB information. Additional ablation studies, which explore various pooling strategies and evaluate performance across all classes, are provided in the supplementary material.
Feature Learning. We compare our dense point sampling based primitive feature learning method with the approach in SymPoint, which uses handcrafted features to represent each individual primitive. SymPoint employs Point Transformer(PTv1) [39] as its backbone network and incorporates hierarchical multi-resolution primitive features to leverage intermediate features, enhancing the decoder’s performance. To evaluate the impact of different point cloud analysis methods, we first replace SymPoint’s backbone with PTv3 [35], maintaining the same decoder architecture as ours to ensure a fair comparison of feature learning techniques in panoptic symbol spotting. As shown in Table.6, with the same decoder configuration, simply switching to PTv3 results in a 1.8% improvement in PQ. When we apply our proposed primitive feature learning method, performance further increases by 8.7% in PQ, demonstrating the significant performance gains achieved by our dense point sampling based feature learning approach.
RGB Color Information. We compare the performance of CADSpotting through the inclusion versus exclusion of color information in the FloorPlanCAD dataset. Table.7 shows that adding RGB information improves the PQ metric by 3.1%.
5 Conclusion
We have introduced CADSpotting, a novel method for panoptic symbol spotting in large-scale CAD drawings. Our approach uses dense point sampling within a unified point cloud framework to represent CAD primitives, combined with a Sliding Window Aggregation technique that integrates weighted voting and Non-Maximum Suppression for precise segmentation. We also present LS-CAD, a dataset of 50 annotated floorplans from diverse building types. Experiments on LS-CAD and FloorPlanCAD datasets have confirmed the effectiveness and scalability of our method, with successful applications in automated 3D reconstruction demonstrating its practical value.
@@ 5 Conclusion
| CADSpotting | FloorPlanCAD | 88.9 | 95.6 | 93.0 | 42.8 | 94.8 | 45.2 |
| CADSpotting | Cross-dataset | 87.3 | 95.4 | 91.4 | 69.9 | 91.6 | 76.3 |
| CADSpotting(SWA) | FloorPlanCAD | / | 53.2 | 93.8 | 56.6 | ||
| CADSpotting(SWA) | Cross-dataset | / | 75.5 | 84.9 | 88.9 | ||
6 More Details on the LS-CAD Dataset
@@
To demonstrate the generalization capability of our model and highlight the value of the LS-CAD dataset, we implement a cross-dataset joint training paradigm. We randomly partition complete large-scale CAD drawings into three subsets: 80% for training, 10% for validation, and 10% for testing. The division of the LS-CAD dataset strictly adheres to the standardized data splitting protocol established in FloorPlanCAD. The training and validation sets from both datasets are subsequently integrated to construct a cross-dataset collaborative training environment. As shown in Table.8, the cross-dataset joint training strategy significantly improves model performanc on the LS-CAD test set, although with slight performance degradation on the original FloorPlanCAD test set. This demonstrates the necessity of the proposed LS-CAD dataset for provding more diverse CAD representations. Notably, CADSpotting+SWA method reaches the best performance of 75.5 on the PQ metric. After cross-dataset joint training, the PQ score for CADSpotting decreases by only 1.2% on FloorPlanCAD, whereas it drops by 4.3% on SymPoint. This indicates that our model exhibits stronger generalization capabilities.
7 Quantitative Evaluation
@@
We further evaluate the necessity and impact of different pooling methods on our approach, using a baseline configuration without a pooling layer. As shown in Table.9, max pooling improves performance by 11.5%, average pooling by 13.0%, and mixed pooling by 14.3% in PQ. Based on these results, we select mixed pooling for our CADSpotting approach.
8 Qualitative Evaluation
@@
9 Sliding Window Implementation Details
The Algorithm.1 details the implementation process of our SWA. In the matrix-NMS [33] process for sliding window aggregation, computing the Intersection over Union (IoU) between all pairs of instance proposal masks is essential for constructing the IoU matrix. However, the number of instance proposals grows substantially with the size of the input drawing. If a dense matrix representation is used, both the time and memory complexities of this operation become prohibitively high.
@@ 9 Sliding Window Implementation Details
10 Automated 3D Interior Reconstruction
@@
| w/o Pooling | 74.6 | 89.3 | 83.6 |
| Max | 86.1 | 94.1 | 91.5 |
| Average | 87.6 | 94.5 | 92.7 |
| Mixed (Max+Average) | 88.9 | 95.6 | 93.0 |
@@
| single door | 91.59 | 97.08 | 94.34 | 75.67 | 90.68 | 83.44 | 87.23 | 92.68 | 94.12 | 87.46 | 92.85 | 94.20 |
| double door | 93.51 | 96.93 | 96.47 | 75.14 | 88.76 | 84.65 | 87.30 | 91.85 | 95.04 | 88.27 | 92.73 | 95.18 |
| sliding door | 96.54 | 98.78 | 97.73 | 83.55 | 50.31 | 83.55 | 94.51 | 96.15 | 98.29 | 91.32 | 95.65 | 95.46 |
| folding door | 57.67 | 87.88 | 65.62 | 39.66 | 95.83 | 41.38 | 70.88 | 83.29 | 85.11 | 52.03 | 82.64 | 62.96 |
@@
| window | 89.60 | 97.24 | 92.14 | 75.49 | 91.87 | 82.17 | 77.13 | 85.24 | 90.49 | 74.53 | 85.06 | 87.62 |
| bay window | 48.15 | 95.26 | 50.55 | 41.92 | 88.51 | 47.37 | 22.25 | 77.38 | 28.75 | 12.00 | 68.83 | 17.44 |
| blind window | 90.57 | 98.17 | 92.25 | 77.71 | 96.14 | 80.83 | 79.56 | 84.85 | 93.76 | 77.28 | 83.91 | 92.10 |
| opening symbol | 46.64 | 76.82 | 60.71 | 37.97 | 72.60 | 52.30 | 36.47 | 79.11 | 46.10 | 20.10 | 72.22 | 27.84 |
| sofa | 88.53 | 96.29 | 91.93 | 71.82 | 82.05 | 87.53 | 76.98 | 90.28 | 85.28 | 73.74 | 87.40 | 84.37 |
| bed | 89.79 | 92.47 | 97.11 | 78.30 | 83.76 | 93.47 | 76.87 | 86.26 | 89.11 | 76.51 | 84.44 | 90.61 |
| chair | 83.58 | 94.04 | 88.88 | 70.05 | 85.54 | 81.89 | 84.42 | 92.26 | 91.51 | 80.85 | 91.71 | 88.16 |
| table | 75.39 | 88.46 | 85.22 | 53.18 | 78.36 | 67.87 | 68.03 | 86.46 | 78.68 | 61.99 | 86.43 | 71.72 |
| TV cabinet | 95.23 | 97.57 | 97.60 | 82.85 | 86.79 | 95.46 | 89.99 | 93.41 | 96.33 | 79.94 | 85.50 | 93.49 |
| Wardrobe | 93.76 | 97.23 | 96.43 | 84.53 | 89.58 | 94.36 | 85.37 | 88.77 | 96.17 | 82.50 | 85.94 | 95.99 |
| cabinet | 81.40 | 91.16 | 89.28 | 65.83 | 81.45 | 80.82 | 71.37 | 82.66 | 86.33 | 69.15 | 83.42 | 82.90 |
| gas stove | 97.34 | 99.17 | 98.16 | 90.08 | 96.15 | 93.69 | 97.10 | 97.89 | 99.19 | 96.66 | 97.61 | 99.03 |
| sink | 87.59 | 95.21 | 92.00 | 76.13 | 91.10 | 83.57 | 84.88 | 90.98 | 93.30 | 83.42 | 90.03 | 92.65 |
| refrigerator | 91.96 | 96.24 | 95.55 | 71.86 | 81.87 | 87.78 | 81.70 | 87.44 | 93.44 | 80.28 | 85.58 | 93.80 |
| air conditioner | 86.53 | 99.02 | 87.39 | 78.66 | 96.36 | 81.64 | 80.59 | 92.04 | 87.56 | 75.45 | 91.70 | 82.27 |
| bath | 79.30 | 93.31 | 84.98 | 63.84 | 83.21 | 76.73 | 66.62 | 79.71 | 83.58 | 62.06 | 77.36 | 80.22 |
| bath tub | 85.00 | 91.92 | 92.47 | 70.96 | 82.90 | 85.61 | 69.94 | 80.14 | 87.27 | 64.82 | 78.13 | 82.96 |
| washing machine | 89.17 | 97.74 | 91.23 | 78.29 | 90.75 | 86.27 | 86.40 | 91.21 | 94.73 | 77.97 | 85.26 | 91.45 |
| squat toilet | 95.35 | 98.25 | 97.05 | 78.70 | 94.34 | 83.42 | 92.58 | 94.79 | 97.67 | 89.99 | 92.68 | 97.09 |
| urinal | 94.32 | 98.76 | 95.51 | 80.33 | 92.08 | 87.24 | 91.80 | 94.84 | 96.79 | 91.31 | 94.79 | 96.32 |
| toilet | 95.21 | 98.08 | 97.07 | 81.39 | 91.92 | 88.55 | 90.81 | 93.01 | 97.63 | 90.48 | 93.53 | 96.74 |
| stairs | 86.47 | 94.28 | 91.72 | 74.68 | 89.68 | 83.27 | 68.56 | 82.02 | 83.59 | 65.50 | 79.52 | 82.36 |
| elevator | 96.15 | 97.53 | 98.58 | 79.08 | 94.76 | 83.44 | 84.44 | 90.52 | 93.28 | 80.64 | 88.59 | 91.03 |
| escalator | 73.73 | 88.37 | 83.44 | 67.98 | 79.15 | 85.88 | 29.80 | 72.72 | 40.98 | 44.64 | 73.70 | 60.56 |
| row chairs | 85.16 | 94.94 | 89.71 | 82.45 | 92.43 | 89.21 | 85.69 | 94.53 | 90.65 | 84.45 | 93.99 | 89.86 |
| parking spot | 89.68 | 95.20 | 94.20 | 84.34 | 92.67 | 91.01 | 72.22 | 80.84 | 89.34 | 72.64 | 82.77 | 87.76 |
| wall | 82.37 | 88.83 | 92.73 | 75.08 | 81.61 | 66.81 | 81.86 | 65.29 | 68.74 | 44.48 | 65.44 | 67.97 |
| curtain wall | 64.80 | 89.88 | 72.09 | 64.17 | 84.45 | 75.98 | 60.25 | 87.36 | 73.90 | 37.62 | 73.21 | 51.39 |
| railing | 75.67 | 92.22 | 82.06 | 62.87 | 85.69 | 73.37 | 37.82 | 76.42 | 49.49 | 28.53 | 70.73 | 40.34 |
| total | 88.91 | 95.65 | 92.96 | 74.59 | 89.26 | 83.56 | 80.24 | 90.17 | 88.98 | 78.36 | 88.59 | 88.46 |

### v4

@@ Introduction
To demonstrate the effectiveness of our approach, we compare it against SOTA methods on the FloorPlanCAD (Fan et al. 2021) and LS-CAD datasets. A detailed comparison of semantic, instance, and panoptic symbol spotting tasks is provided in the following paragraphs.
@@
Semantic Symbol Spotting As shown in Table.2, we compare our method with PanCADNet (Fan et al. 2021), CADTransformer (Fan et al. 2022), GAT-CADNet (Zheng et al. 2022), and SymPoint (LIU et al. 2024). Our CADSpotting achieves the highest performance in both the F1 and wF1 metrics.
@@
Instance Symbol Spotting We also evaluate our method against DINO (Zhang et al. 2022) and SymPoint for instance symbol spotting, using the bounding box maximization of predicted mask to compute the AP metric, consistent with SymPoint. As shown in Table.3, our method achieves superior results , with improvements of 4.3 p.p. AP50, 10.9 p.p. AP75 and 13.5 p.p. mAP over SymPoint.
@@
Panoptic Symbol Spotting We compare our method with the same methods used in the semantic symbol spotting comparison. The results, presented in Table.4, show that our approach surpasses SymPoint and other prior methods across all metrics, including Total/Thing/Stuff PQ, SQ and RQ. Additionally, our approach offers superior generalization across diverse CAD primitives, thanks to its dense point sampling based feature representation.
| Method | Panoptic | Semantic | Instance | ||
| PQ | SQ | RQ | F1 | mAP | |
| SPv2 | 90.1 | 96.3 | 93.6 | 89.5 | 60.1 |
| Ours | 87.4 | 93.5 | 93.4 | 92.8 | 66.3 |
| Method | Panoptic | Semantic | Instance | ||
| PQ | SQ | RQ | F1 | mAP | |
| SPv2 + BP | 26.7 | 82.0 | 32.6 | 55.2 | 15.1 |
| Ours + BP | 58.8 | 83.7 | 70.2 | 88.7 | 48.3 |
| Ours + SWA | 75.5 | 80.9 | 93.3 | 93.5 | 57.5 |
We then compare our method with the concurrent work SymPoint-V2 (SPv2) (Liu et al. 2024), which incorporates primitive types into feature vectors. Our first comparison is conducted on the FloorPlanCAD dataset. As shown in Table.5, although SPv2 achieves up to 2.7 percentage points higher PQ in panoptic segmentation, our method delivers a 3.3-point gain in F1 score and a 6.2-point improvement in mAP. While SPv2 performs well on FloorPlanCAD due to its reliance on four fixed primitive types, this limits its generalization to more diverse CAD symbol sets. In contrast, our method relies solely on coordinate attributes, enabling more flexible and robust feature learning. Furthermore, SPv2 depends on CAD layer information, which is often inconsistent or unavailable in real-world scenarios. Table. 6 further supports this analysis. On the LS-CAD dataset, where layer annotations are unavailable, our method significantly outperforms SPv2 across all metrics, demonstrating stronger robustness in practical applications.
We also evaluate the effectiveness of Sliding Window Aggregation (SWA) versus Block Partitioning (BP) in CADSpotting on the LS-CAD dataset. In the BP setting, we divide large-scale CAD drawings into non-overlapping blocks matching the dimensions used in FloorPlanCAD. To ensure unique primitive representation, only those with starting points inside a block are retained. As shown in Table 6, SWA substantially outperforms BP in panoptic quality (PQ), achieving a 16.7-point improvement, with only a modest 2.8-point reduction in segmentation quality (SQ). These results demonstrate that CADSpotting, when combined with SWA, achieves accurate and scalable symbol detection in large-scale CAD drawings.
Qualitative Comparison
@@
| Ours | PTv3 | 87.4 | 93.5 | 93.4 |
Ablation Studies
We further present ablation studies on feature learning. Additional ablation studies, which explore various pooling strategies and the use of color information, and evaluate performance across all classes, are provided in the supplementary material.
Feature Learning We compare our dense point sampling based primitive feature learning method with the approach in SymPoint, which uses handcrafted features to represent each individual primitive. SymPoint employs Point Transformer(PTv1) (Zhao et al. 2021) as its backbone network and incorporates hierarchical multi-resolution primitive features to leverage intermediate features, enhancing the decoder’s performance. To evaluate the impact of different point cloud analysis methods, we first replace SymPoint’s backbone with PTv3 (Wu et al. 2024), maintaining the same decoder architecture as ours to ensure a fair comparison of feature learning techniques in panoptic symbol spotting. As shown in Table.7, with the same decoder configuration, simply switching to PTv3 results in a 1.8 p.p. improvement in PQ. When we apply our proposed primitive feature learning method, performance further increases 7.2 p.p. in PQ, demonstrating the significant performance gains achieved by our dense point sampling based feature learning approach.
Conclusion
We have introduced CADSpotting, a novel method for panoptic symbol spotting in large-scale CAD drawings. Our approach uses dense point sampling within a unified point cloud framework to represent CAD primitives, combined with a Sliding Window Aggregation technique that integrates weighted voting and Non-Maximum Suppression for precise segmentation. We also present LS-CAD, a dataset of 45 annotated floorplans from diverse building types. Experiments on LS-CAD and FloorPlanCAD datasets have confirmed the effectiveness and scalability of our method, with successful applications in automated 3D reconstruction demonstrating its practical value.
@@ Conclusion
| CADSpotting | FloorPlanCAD | 87.4 | 93.5 | 93.4 | 52.9 | 90.4 | 58.5 |
| CADSpotting | Cross-dataset | 84.0 | 91.3 | 91.9 | 58.8 | 83.7 | 70.2 |
| CADSpotting(SWA) | FloorPlanCAD | / | 60.1 | 90.9 | 66.1 | ||
| CADSpotting(SWA) | Cross-dataset | / | 75.5 | 80.9 | 93.3 | ||
More Details on the LS-CAD Dataset
@@
To demonstrate the generalization capability of our model and highlight the value of the LS-CAD dataset, we implement a cross-dataset joint training paradigm. We randomly partition complete large-scale CAD drawings into three subsets: 80% for training, 10% for validation, and 10% for testing. The division of the LS-CAD dataset strictly adheres to the standardized data splitting protocol established in FloorPlanCAD. The training and validation sets from both datasets are subsequently integrated to construct a cross-dataset collaborative training environment. As shown in Table.8, the cross-dataset joint training strategy significantly improves model performanc on the LS-CAD test set, although with slight performance degradation on the original FloorPlanCAD test set. This demonstrates the necessity of the proposed LS-CAD dataset for provding more diverse CAD representations. Notably, CADSpotting+SWA method reaches the best performance of 75.5 on the PQ metric. After cross-dataset joint training, the PQ score for CADSpotting decreases by only 3.4 p.p. on FloorPlanCAD, whereas it drops by 4.3 p.p. on SymPoint. This indicates that our model exhibits stronger generalization capabilities.
Quantitative Evaluation
@@
We further evaluate the necessity and impact of different pooling methods on our approach, using a baseline configuration without a pooling layer. As shown in Table.9, max pooling improves performance by 5.0 p.p., average pooling by 5.1 p.p., and mixed pooling by 5.5. p.p. in PQ. Based on these results, we select mixed pooling for our CADSpotting approach.
Ablation Study on Prior Color Information.
To evaluate the necessity of color prior, we incorporate RGB channels as an optional input. As shown in Table 10, adding color improves PQ by 1.5 p.p. and SQ by 2.1 p.p., while slightly decreasing RQ by 0.4 p.p. More importantly, the geometry-based representation (w/o Color) achieves strong baseline performance, demonstrating that our model maintains high recognition quality even without color information.
Qualitative Evaluation
@@
Sliding Window Implementation Details
The Algorithm.1 details the implementation process of our SWA. In the matrix-NMS (Wang et al. 2020) process for sliding window aggregation, computing the Intersection over Union (IoU) between all pairs of instance proposal masks is essential for constructing the IoU matrix. However, the number of instance proposals grows substantially with the size of the input drawing. If a dense matrix representation is used, both the time and memory complexities of this operation become prohibitively high.
@@
Automated 3D Interior Reconstruction
@@
| w/o Pooling | 81.9 | 89.4 | 91.6 |
| Max | 86.9 | 93.3 | 93.2 |
| Average | 87.0 | 93.1 | 93.4 |
| Mixed (Max+Average) | 87.4 | 93.5 | 93.4 |
| Method | PQ | SQ | RQ |
| w/ color | 87.4 | 93.5 | 93.3 |
| w/o color | 88.9 | 95.6 | 93.0 |
@@
| single door | 91.90 | 96.33 | 95.40 | 84.11 | 91.34 | 92.08 | 87.23 | 92.68 | 94.12 | 87.46 | 92.85 | 94.20 |
| double door | 94.10 | 96.46 | 97.55 | 87.23 | 91.70 | 95.13 | 87.30 | 91.85 | 95.04 | 88.27 | 92.73 | 95.18 |
| sliding door | 96.50 | 98.02 | 98.46 | 92.88 | 95.04 | 97.73 | 94.51 | 96.15 | 98.29 | 91.32 | 95.65 | 95.46 |
| folding door | 77.06 | 89.23 | 86.36 | 73.03 | 92.36 | 79.07 | 70.88 | 83.29 | 85.11 | 52.03 | 82.64 | 62.96 |
@@
| window | 85.41 | 91.06 | 93.80 | 77.71 | 84.43 | 92.04 | 77.13 | 85.24 | 90.49 | 74.53 | 85.06 | 87.62 |
| bay window | 51.18 | 90.19 | 56.74 | 43.23 | 91.50 | 47.25 | 22.25 | 77.38 | 28.75 | 12.00 | 68.83 | 17.44 |
| blind window | 86.00 | 92.25 | 93.22 | 77.99 | 86.97 | 89.68 | 79.56 | 84.85 | 93.76 | 77.28 | 83.91 | 92.10 |
| opening symbol | 52.67 | 77.06 | 68.35 | 34.51 | 74.91 | 46.07 | 36.47 | 79.11 | 46.10 | 20.10 | 72.22 | 27.84 |
| sofa | 89.68 | 96.78 | 92.67 | 85.50 | 92.42 | 92.51 | 76.98 | 90.28 | 85.28 | 73.74 | 87.40 | 84.37 |
| bed | 82.12 | 88.75 | 92.54 | 79.51 | 86.32 | 92.11 | 76.87 | 86.26 | 89.11 | 76.51 | 84.44 | 90.61 |
| chair | 84.72 | 93.66 | 90.46 | 86.51 | 94.63 | 91.42 | 84.42 | 92.26 | 91.51 | 80.85 | 91.71 | 88.16 |
| table | 77.42 | 89.58 | 86.42 | 73.30 | 87.39 | 83.87 | 68.03 | 86.46 | 78.68 | 61.99 | 86.43 | 71.72 |
| TV cabinet | 94.56 | 97.08 | 97.40 | 83.23 | 85.80 | 97.01 | 89.99 | 93.41 | 96.33 | 79.94 | 85.50 | 93.49 |
| Wardrobe | 91.56 | 94.54 | 96.85 | 88.97 | 92.63 | 96.05 | 85.37 | 88.77 | 96.17 | 82.50 | 85.94 | 95.99 |
| cabinet | 80.10 | 89.04 | 89.95 | 73.11 | 84.25 | 86.78 | 71.37 | 82.66 | 86.33 | 69.15 | 83.42 | 82.90 |
| gas stove | 97.89 | 99.21 | 98.67 | 96.76 | 98.39 | 98.35 | 97.10 | 97.89 | 99.19 | 96.66 | 97.61 | 99.03 |
| sink | 87.08 | 93.71 | 92.93 | 84.87 | 91.49 | 92.77 | 84.88 | 90.98 | 93.30 | 83.42 | 90.03 | 92.65 |
| refrigerator | 91.05 | 95.32 | 95.52 | 80.35 | 84.60 | 94.97 | 81.70 | 87.44 | 93.44 | 80.28 | 85.58 | 93.80 |
| air conditioner | 87.43 | 98.17 | 89.06 | 82.21 | 95.67 | 85.93 | 80.59 | 92.04 | 87.56 | 75.45 | 91.70 | 82.27 |
| bath | 75.50 | 88.30 | 85.50 | 69.96 | 81.78 | 85.54 | 66.62 | 79.71 | 83.58 | 62.06 | 77.36 | 80.22 |
| bath tub | 83.67 | 88.14 | 94.92 | 76.50 | 81.37 | 94.01 | 69.94 | 80.14 | 87.27 | 64.82 | 78.13 | 82.96 |
| washing machine | 88.26 | 95.38 | 92.54 | 82.34 | 88.59 | 92.95 | 86.40 | 91.21 | 94.73 | 77.97 | 85.26 | 91.45 |
| squat toilet | 92.85 | 96.01 | 96.71 | 92.41 | 96.23 | 96.03 | 92.58 | 94.79 | 97.67 | 89.99 | 92.68 | 97.09 |
| urinal | 94.46 | 97.76 | 96.62 | 91.24 | 94.17 | 96.90 | 91.80 | 94.84 | 96.79 | 91.31 | 94.79 | 96.32 |
| toilet | 95.83 | 97.70 | 98.09 | 91.22 | 92.96 | 98.13 | 90.81 | 93.01 | 97.63 | 90.48 | 93.53 | 96.74 |
| stairs | 84.41 | 92.37 | 91.38 | 80.19 | 88.51 | 90.60 | 68.56 | 82.02 | 83.59 | 65.50 | 79.52 | 82.36 |
| elevator | 95.06 | 97.57 | 97.43 | 93.26 | 96.46 | 96.69 | 84.44 | 90.52 | 93.28 | 80.64 | 88.59 | 91.03 |
| escalator | 74.96 | 87.62 | 85.55 | 70.43 | 85.84 | 82.05 | 29.80 | 72.72 | 40.98 | 44.64 | 73.70 | 60.56 |
| row chairs | 85.73 | 94.58 | 90.65 | 84.15 | 93.96 | 89.55 | 85.69 | 94.53 | 90.65 | 84.45 | 93.99 | 89.86 |
| parking spot | 88.54 | 93.15 | 95.05 | 84.48 | 90.24 | 93.62 | 72.22 | 80.84 | 89.34 | 72.64 | 82.77 | 87.76 |
| wall | 73.21 | 81.42 | 89.91 | 62.55 | 73.32 | 85.31 | 44.88 | 65.29 | 68.74 | 44.48 | 65.44 | 67.97 |
| curtain wall | 58.66 | 82.61 | 71.01 | 50.49 | 78.00 | 64.73 | 60.25 | 87.36 | 73.90 | 37.62 | 73.21 | 51.39 |
| railing | 67.25 | 88.38 | 76.10 | 59.50 | 83.88 | 70.94 | 37.82 | 76.42 | 49.49 | 28.53 | 70.73 | 40.34 |
| total | 87.36 | 93.40 | 93.54 | 81.89 | 89.42 | 91.58 | 80.24 | 90.17 | 88.98 | 78.36 | 88.59 | 88.46 |